OpenStudy (anonymous):

Given the two expressions shown below: posted below Which statement best describes the two expressions? Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational

OpenStudy (anonymous):

\[\sqrt{3} + \sqrt{2}\] \[\sqrt{3}+\sqrt{6}\]

OpenStudy (anonymous):

@kohai

OpenStudy (kohai):

Rational numbers are numbers that either repeat or terminate. So for example, 0.25, sqrt16, etc. Irrational numbers are numbers that do not repeat or terminate. For example, pi, e, sqrt2, etc

OpenStudy (kohai):

So basically, if it doesn't eventually turn into a whole number, it's going to be irrational. So in this case, both of them are going to be irrational expressions

OpenStudy (anonymous):

thanks posting next lol

OpenStudy (kohai):

Please metal and fan me :) It makes it so that way I can gradually increase my smart score!